OpenStudy (anonymous):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x - 12)^2 + 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

I know it's up three

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (anonymous):

But i get confused with left and right

OpenStudy (anonymous):

Is left positive?

jimthompson5910 (jim_thompson5910):

since you have x-12, you go to the right 12 units in general, you have x-h and the h is how you move horizontally

jimthompson5910 (jim_thompson5910):

if h > 0, then you move to the right if h < 0, then you move to the left

OpenStudy (anonymous):

so right 12 and up 3?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

wait but since h is -12 and it's lower than 0, would it be left 12?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it's x - h so you replace just the h with 12 to get x - 12 that means h = 12

jimthompson5910 (jim_thompson5910):

if you had x+12, then x+12 = x - (-12) compare that to x - h and we see that h = -12 now

OpenStudy (anonymous):

ooooh gotcha